# Soft-QMIX: Integrating Maximum Entropy For Monotonic Value Function Factorization

### v2

@@ Appendix C Joint Policy Iteration
Let $\pi_{jt}^{i}$ represent the policy after the $i$-th update. By Lemma B.2, the sequence $Q^{\pi_{jt}^{i}}$ is monotonically increasing. Given that $Q^{\pi}$ is bounded above for $\pi_{jt}\in\Pi$ due to bounded rewards and entropy terms, the sequence converges to some $\pi^{*}_{jt}$. It remains to demonstrate that the policy $\pi^{*}_{jt}$ is indeed the optimal policy. At convergence, we infer that $J_{\pi^{*}_{jt}}(s)\geq J_{\pi_{jt}}(s)$ for all $\pi_{jt}\in\Pi$, if $\pi_{jt}\neq\pi^{*}_{jt}$. Using the same iterative argument as in the proof of Lemma B.2, we get $Q^{\pi^{*}_{jt}}(s,\mathbf{u})\geq Q^{\pi}(s,a)$ for all $(s,\mathbf{u})\in S\times\mathbf{U}$, that is, the soft value of any other policy in $\Pi$ is lower than that of the converged policy. Hence $\pi^{*}_{jt}$ is optimal in $\Pi$. ∎
Appendix D TD($\lambda$)
In this paragraph, we will derive the expression of TD($\lambda$) in the context of Maximum Entropy Reinforcement Learning. Specifically, we aim to demonstrate the derivation of the following equation:
Proof.
The one-step TD error can be expressed in the following form:
$\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$ (25)
Next, the k-step TD error can be represented in the following form:
$\epsilon_{t}^{(1)}:=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})=\delta^{V}_{t}$ (26)
$\epsilon_{t}^{(2)}:=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma r_{t+1}{\color[rgb]{1,0,0}-\log\pi_{jt}(\mathbf{u_{t+1}}|s_{t+1})}+\gamma^{2}V_{tot}(s_{t+1})=\delta^{V}_{t}+\gamma\delta^{V}_{t+1}$
$\epsilon_{t}^{(k)}=\sum_{l=0}^{k-1}\gamma^{l}\delta^{V}_{t+l}$
$\epsilon_{t}^{td(\lambda)}:=(1-\lambda)\left(\epsilon_{t}^{(1)}+\lambda\epsilon_{t}^{(2)}+\lambda^{2}\epsilon_{t}^{(3)}+\cdots\right)$ (27)
$=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{V}$
Appendix E Soft-Q Learning and Policy Gradient
Theorem E.1.
If $Q$-function is parameterized as $Q_{\theta}$, and $V_{\theta},\pi_{\theta}$ are defined as below:
$V_{\theta}(s)$ :-$\coloneq\alpha\log\mathbb{E}_{\mathbf{u}\sim\pi_{jt}}\left[\exp\left(Q_{\theta}(s,\mathbf{u})/\alpha\right)\right],:$ (28)
$\pi_{jt_{\theta}}(\mathbf{u}|s)$ :-$\coloneq\exp\left(Q_{\theta}(s,\mathbf{u})-V_{\theta}(s)\right)/\alpha.:$
and $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right].$ Then, we have
$\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}$ $\left[\|Q_{\theta}(s,\mathbf{u})-y\|^{2}\right]$ (29)
$=\alpha\nabla_{\theta}D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$
$+\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\|V_{\theta}(s)-y\|^{2}\right]$
Proof.
The proof is heavily based on [25]. From assumptions described in Eq. 11 we have:
$Q_{\theta}(s,\mathbf{u})=V_{\theta}(s)+\alpha\log\pi_{jt_{\theta}}(\mathbf{u}|s)$ (30)
For simplicity, we further define notation $\Delta_{t}$ as below.
$\Delta_{t}$ $=\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})-Q_{\theta}(s_{t},\mathbf{u_{t}}).$
From assumption, we have $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right]$. Thus, the relationship between $\Delta_{t}$ and $y_{t}$ can be writen as:
$\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[y_{t}]$ $=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right]$ (31)
$=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[Q_{\theta}(s_{t},\mathbf{u_{t}})+\Delta_{t}]$
$=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[V_{\theta}(s_{t})+\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]+\Delta_{t}].$
Here are two equations that will be used in the following proof:
$\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s)\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u}|s)\big{)}\big{]}$ (32)
$=$ $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s)\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u}|s)]\big{)}\big{]}.$
$\mathbb{E}_{a\sim\pi}\big{[}\nabla\log\pi(a)f(s)\big{]}=f(s)\sum_{a}\pi(a)\frac{\nabla\pi(a)}{\pi(a)}$ (33)
$=f(s)\sum_{a}\nabla\pi(a)=f(s)\nabla\sum_{a}\pi(a)=0,$
where $f$ can be any function that does not conditioned on $a$, and $\sum_{a}\pi(a)=1$. Let’s consider the gradient of the $Q$-function’s MSE loss:
$\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\frac{1}{2}\left(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t}\right)^{2}\right]\bigg{|}_{\pi=\pi_{\theta}}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t})\right]\bigg{|}_{\pi=\pi_{\theta}}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\left(V_{\theta}(s_{t})+\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-y_{t}\right)\right]$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\big{(}V_{\theta}(s_{t})+\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-V_{\theta}(s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s_{t})+\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u}|s)\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}$
$\qquad-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}$
$\qquad-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-$
$\qquad\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}.$
for simplicity, we ignore the $|_{\pi=\pi_{\theta}}$ after the first two lines of equations. The second equation use Eq. 30. The third equation use Eq. 31. The fourth equation cancled out $V_{\theta}(s_{t})$. The fifth equation use Eq. 30 again. The sixth equation expand out terms. The seventh equation use Eq. 32. The last equation use Eq. 33.
Let’s consider the gradient of the KL divergence.
$\nabla_{\theta}\alpha D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$
$=\nabla_{\theta}\mathbb{E}_{\pi_{jt}}\big{[}\alpha\log\pi_{jt}(\mathbf{u}|s)-y(\mathbf{u},s)+\alpha\log Z(s)\big{]}$
$=\nabla_{\theta}\mathbb{E}_{\pi_{jt}}\big{[}\alpha\log\pi_{jt}(\mathbf{u}|s)-y(\mathbf{u},s)\big{]}$
$=\nabla_{\theta}[\alpha\sum_{\mathbf{u}}\pi_{jt}(\mathbf{u}|s)\log\pi_{jt}(\mathbf{u}|s)]-\nabla_{\theta}[\sum_{\mathbf{u}}\pi_{jt}(\mathbf{u}|s)y(\mathbf{u},s)]$
$=\alpha\sum_{\mathbf{u}}\nabla_{\theta}\pi_{jt}(\mathbf{u}|s)\log\pi_{jt}(\mathbf{u}|s)-\sum_{\mathbf{u}}\nabla_{\theta}\pi_{jt}(\mathbf{u}|s)y(\mathbf{u},s)$
$=\mathbb{E}_{\pi_{jt}}\big{[}\alpha\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\log\pi_{jt}(\mathbf{u}|s)-\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}y(\mathbf{u},s)\big{]}.$
Note that $V_{\theta}(s)+\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u}|s)]$ is not a function of $\mathbf{u}$. With Eq. 33, we have:
$\nabla_{\theta}\alpha D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$ (34)
$=\mathbb{E}_{\pi_{jt}}\big{[}\alpha\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\log\pi_{jt}(\mathbf{u}|s)-\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\Delta_{t}\big{]}$
Finally, we got:
$\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\frac{1}{2}\left(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t}\right)^{2}\right]\bigg{|}_{\pi=\pi_{\theta}}$
$=\nabla_{\theta}\alpha^{2}D_{KL}\big{(}\pi_{jt}(\cdot|s_{t})\|\frac{exp(y(s)/\alpha)}{Z(s)}\big{)}+\frac{1}{2}\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}(V_{\theta}(s_{t})-y_{t})^{2}\big{]},$
where $y(s)$ is the distribution of $y_{t}$ at state $s$.
Appendix F Unbiased Assumption
We assume that $\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u_{t}})$ is an unbiased estimate of $E_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}\sum_{i}f_{i}(Q_{i}(s_{t},\mathbf{u_{t}}))\right]$ in the main paper. Practically, this assumption might fail, leading to non-monotonic increases in the policy improvement step. Here, we initially explain the impact of this error and derive an upper bound for it.
Conclusion: the error’s upper bound is proportional to $|Q_{tot}-\sum_{i}f_{i}(Q_{i})|$ and $\sqrt{D_{KL}\big{(}\pi_{jt}^{\text{old}}|\pi_{jt}^{\text{new}}\big{)}}.$ We use Equation 14 as the loss function, ensuring that the first term decreases throughout training. Additionally, by selecting a sufficiently small learning rate, the second term will also remain minimal.
Derivative: We can replace both $Q_{\theta}$ and $y_{t}$ in Theorem 4.2 with $\sum_{i}f_{i}(Q_{i})$, and get that minimizing the L2 norm described in Equation 10 is equivalent to minimizing the td loss of value function and minimizing the KL divergence described as follow,
$\pi_{jt}^{\text{new}}=\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{arg\,min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(\sum_{i}f_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (35)
[…]
Similar to Equation 22, during policy improvement step, we have:
$\mathbb{E}_{\mathbf{u}_{t}\sim\pi_{jt}^{\text{new}}}\left[\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t},u^{i}_{t}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t}\mid s_{t})\right]\leq V^{\prime\pi_{jt}^{\text{old}}}(s_{t}).$ (36)
During policy evaluation step, we have:
@@ Appendix F Unbiased Assumption
$\leq r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}+\gamma\epsilon_{1}$
@@
where $\epsilon$ is s the cumulative sum of $\epsilon_{1}$ and $\epsilon_{2}$ over time.$\epsilon_{1}$ and $\epsilon_{2}$ are defined as follows:
@@
$\epsilon_{2}$ $=\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
@@
$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
$=\mathbb{E}_{s_{t+1}\sim p}\left[\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{old}}}\big{[}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{]}\right]$
$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
$=\mathbb{E}_{s_{t+1}\sim p}\big{[}\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{(}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{)}\big{]}$
$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{]},$
where $C=\underset{\mathbf{u}_{t+1}}{\mathrm{arg\,max}}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality [21], we have:
$|\epsilon_{1}+\epsilon_{2}|$ $\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big{]}$ (40)
$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sqrt{2D_{KL}\big{(}\pi_{jt}^{\text{old}}(\cdot|s_{t+1})|\pi_{jt}^{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$
Appendix G Hyper-network Structure
@@
Figure 4: (a) The overall network architecture. Agent 1 to Agent n takes local observations as input and outputs a distribution over local actions $u^{i}$. Functions $g_{i}$ and $f_{i}$ are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values. Both the input and output of the $g_{i}$ and $f_{i}$ networks are k-dimensional, where $k$ is the action dimension. The output of the $f_{i}$ function, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled. The Mixer network takes local $Q_{i}$ as input and outputs the global $Q_{tot}$. Green blocks utilize a recurrent neural network as the backbone, with all agents sharing parameters. For decentralized execution, only the green network is needed, and the maximum values of their output are selected. (b) The blue blocks’ network structure. The red blocks represent the hyper-network, which takes the global state $s_{t}$ as input and outputs the network’s weights and biases. The weights are constrained to be non-negative. The activation function is denoted by $\sigma(\cdot)$. The mixer network takes an n-dimensional Q function as input and outputs a one-dimensional $Q_{tot}$.
Figure 4 shows the overall network structure. Specifically, function $f_{i}$ and function $g_{i}$ share the same network structure as the mixer network, but while the mixer network’s input dimension is (batch_size, agent_num) and its output dimension is (batch_size, 1), both the input and output dimensions of function $f_{i}$ and function $g_{i}$ are (batch_size, 1). Additionally, function $g_{i}$ introduces an additional residual from input to output. The network hyperparameters are as shown in the table below.
layer name hypernet_embed embed_dim num_layer
@@
function $g_{i}$ 64 32 2
function $f_{i}$ 64 - 1
Table 5: Hyper-parameters used for hyper-networks.
Appendix H Comparison with Baselines
The following paragraphs present the differences between our method and those based on the AC:

### v3

@@ Acknowledgment
We define the value function as $V^{\prime\pi_{jt}^{\text{new}}}(s)=\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{new}}}(o^{i}_{t+1},u^{i}_{t+1}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}$. Then, similar to Equation 18, during policy improvement step, we have:
$\mathbb{E}_{\mathbf{u}_{t}\sim\pi_{jt}^{\text{new}}}\left[\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t},u^{i}_{t}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t}\mid s_{t})\right]\geq V^{\prime\pi_{jt}^{\text{old}}}(s_{t}).$ (21)
By expanding the soft Bellman equation, we have:
@@
$\leq r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}+\gamma\epsilon_{1}$
@@
where $p$ is the transition function $\mathcal{P}(s_{t},a_{t},\cdot)$, and $\epsilon$ is the cumulative sum of $\epsilon_{1}$ and $\epsilon_{2}$ over time. $\epsilon_{1}$ and $\epsilon_{2}$ are defined as follows:
@@
$\epsilon_{2}$ $=\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
@@
$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
$=\mathbb{E}_{s_{t+1}\sim p}\left[\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{old}}}\big{[}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{]}\right]$
$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
$=\mathbb{E}_{s_{t+1}\sim p}\big{[}\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{(}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{)}\big{]}$
$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})||Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|\big{]}$
$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big{]},$
where $C=\underset{\mathbf{u}_{t+1}}{\max}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality (Pinsker 1964), we have:
$\epsilon\leq\frac{\gamma}{1-\gamma}|\epsilon_{1}+\epsilon_{2}|$ $\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big{]}$ (25)
$\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sqrt{2D_{KL}\big{(}\pi_{jt}^{\text{old}}(\cdot|s_{t+1})|\pi_{jt}^{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$
Appendix C C. Hyper-network Structure
@@
Figure 7: (a) The overall network architecture. Single Q-Net takes local observations as input and outputs a distribution over local actions $u^{i}$, also denoted as local $Q_{i}$. Functions OPTs are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values. Both the input and output of the OPTs are k-dimensional, where $k$ is the action dimension. The output of the OPT, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled. The Mixer network takes local $Q_{i}$ as input and outputs the global $Q_{tot}$. Green blocks utilize a recurrent neural network as the backbone, with all agents sharing parameters. For decentralized execution, only the green network is needed, and the maximum values of their output are selected. (b) The blue blocks’ network structure. The red blocks represent the hyper-network, which takes the global state $s_{t}$ as input and outputs the network’s weights and biases. The weights are constrained to be non-negative. The activation function is denoted by $\sigma(\cdot)$. The mixer network takes an n-dimensional Q function as input and outputs a one-dimensional $Q_{tot}$.
Figure 7 shows the overall network structure. Specifically, OPTs share the same network structure as the mixer network, but while the mixer network’s input dimension is (batch_size, agent_num) and its output dimension is (batch_size, 1), both the input and output dimensions of OPTs are (batch_size, action_num). The network hyperparameters are as shown in the table below. In the SMAC-v2 experiment, we select a single layer for the OPT model to ensure fairness in comparison. Using two layers of OPT would significantly increase the number of parameters, making it less comparable to the baseline algorithms. While this may slightly reduce our method’s performance, we adopt this choice for a fair comparison.
layer name hyper-net_embed embed_dim num_layer
@@
OPT 64 - 1
Table 5: Hyperparameters used for hyper-networks.
Appendix D D. Comparison with Baselines
The following paragraphs present the differences between our method and those based on the Actor-Critic framework: